0.14941 wt% aqueous KCl solution
at 25°c was electrolyzed in a hittorf apparatus using two Ag-AgCl electrods .The cathode reaction was AgCl+ e ~ Ag + Cl- , the anode reaction was the reverse of this .After the experiment ,it was found that 143.34 C charge had been flown through the apparatus and that the cathode compartment contained 120.99g of solution
that was 0.19404wt% KCl by weight.Calculate t+ and t- in the KCl solution
used in the experiment ,neglect the transport of water by ions.
1
Expert's answer
2019-07-29T04:30:13-0400

Solution.

"(t+) + (t-) = 1"

To determine the number of transfer of potassium ions and chlorine, we write the following formula:


"(t+) = \frac{F \times \Delta n(k)}{Q}""\Delta w = w2-w1"

"\Delta w = \frac{0.19404-0.14941}{100} = 0.04463"

"\Delta w = \frac{m(substance) \times 100}{m(solution)}"

"m(substance) = \frac{\Delta w \times m(solution)}{100}"

m(substance) = 0.0540 g


"n(KCl) = \frac{m}{M}"

n(KCl) = 0.000724 mole


"\Delta n(KCl) = 0.000724 mole""(t+) = \frac{96500 \times 0.000724}{143.34} = 0.49"

"(t-) = 1 - (t+)"

"(t-) = 0.51"

(t+) = 0.49

(t-) = 0.51

Answer:

(t+) = 0.49

(t-) = 0.51